# The properties of GSN 069 accretion disk from a joint X-ray and UV spectral analysis: stress-testing quasi-periodic eruption models

### v1

@@ 2.2 HST/STIS UV Spectra
Although the signal-to-noise (S/N) ratio is high, no strong broad emission lines typically seen in Seyfert galaxies and quasars (Shull et al., 2012; Stevans et al., 2014) are detected. In this seems, the UV spectra is mostly ‘featureless’, in comparison with any standard AGN. This absence of lines is unsurprising, as GSN 069 lacks many of the classic AGN features–not only broad optical emission lines (indicating the absence of a mature broad line region, Wevers et al. 2022), but also no hot dust “torus” emission nor hard X-ray corona emission (Miniutti et al., 2019). A comparison to optically selected TDEs may also be warranted: their broad emission lines, usually, used to tell them apart from other optical transients, are associated with their early time ($\leq$ 1 year) ‘UV/optical excess’ emission, whose origin is still debated but is known not to be power by direct disk emission; these lines and this emission component are not present at their later times ($>$ 1 year) when pure disk emission dominates the multi-wavelength TDE emission (Charalampopoulos et al., 2022; van Velzen et al., 2019; Mummery et al., 2024b). If GSN 069 does indeed have a TDE origin, then this early-time UV/optical emission phase was never observed in GSN 069, and all the data presented in this work were taken many years after this initial phase.
@@ 3.1 Fitting Setup and Modeling in the Newtonian Limit
More physically meaningful than $R_{\rm out}/R_{\rm in}$ and $R_{\rm in}^{*}$ would be estimates of $R_{\rm out}$ (in units of the gravitational radius, $R_{\rm g}$) and $R_{\rm in}$. The former would allow for direct comparison with $R_{\rm orb}$, while the latter, when associated with the innermost stable circular orbit (ISCO), is directly correlated with the black hole mass ($M_{\rm BH}$), and would allow for $M_{\rm BH}$ estimates. However, given that in the Newtonian limit of diskSED $i$ and $a$ cannot be marginalized, $R_{\rm out}/R_{\rm g}$ and $M_{\rm BH}$ estimates can only be obtained by assuming ad hoc values (or probability distributions) of $a$ and $i$ (see e.g., Guolo & Mummery, 2024, Wevers et al., in prep). Instead, given the high-quality available data on GSN 069, i.e., two epochs of very high S/N X-ray spectrum and wide UV coverage, estimates on inclination and spin may be obtainable via introduction of relativistic effects into the disk modeling. This is the focus of the next section.
@@ 3.2 Relativistic Disk Modeling
$M_{\rm BH}\,[M_{\odot}]$ $5.5^{+1.1}_{-0.6}\times 10^{6}$ $1.3^{+0.5}_{-0.3}\times 10^{6}$
@@
$\lambda_{\rm Edd}^{\rm 2014}$ $0.07^{+0.02}_{-0.01}$ $0.31^{+0.09}_{-0.07}$
$\lambda_{\rm Edd}^{\rm 2018}$ $0.05^{+0.01}_{-0.01}$ $0.23^{+0.06}_{-0.05}$
$\lambda_{\rm Edd}^{\rm 2018}/\lambda_{\rm Edd}^{\rm 2014}$ $0.69\pm 0.01$ $0.73\pm 0.01$
@@
Following this approach, we obtain $M_{\rm BH}$ $=5.5^{+1.1}_{-0.6}\times 10^{6}$ $M_{\odot}$ (high spin mode) and $M_{\rm BH}$ $=1.3^{+0.5}_{-0.3}\times 10^{6}$ $M_{\odot}$ (low spin mode). These values are consistent with host-galaxy scaling relations. The nuclear stellar velocity dispersion of GSN 069’s host is $\sigma_{*}=63\pm 4$ km s-1 (Wevers et al., 2022), which translates into $M_{\rm BH}$ estimates of ${\rm few}\times 10^{5}-{\rm few}\times 10^{6}$ $M_{\odot}$ (for the central value only, not adding the intrinsic dispersions) using distinct $M_{\rm BH}$ $-\sigma_{*}$ relations. These relations, however, have systematic $1\sigma$ dispersions of approximately ± ${\pm\,0.5\ \rm dex}$. Thus, even considering the full spin posterior distribution, our $M_{\rm BH}$ estimates are more precise than those derived from host-scaling relations, as our 99% credible interval is (0.8, 8.0)×$\times 10^{6}$ $M_{\odot}$, more importantly, our solutions for distinct black hole masses have distinct associated spins and inclination distributions (Table 2).
@@
With the multi-wavelength fitting performed, we can now obtain more precise estimates of the bolometric disk luminosity ($L_{\rm Bol}$) compared to estimates based solely on X-ray data. For the two epochs, we find $L_{\rm Bol}^{2014}=5.1^{+0.2}_{-0.1}\times 10^{43}$ erg s-1, and $L_{\rm Bol}^{2018}=3.7^{+0.2}_{-0.1}\times 10^{43}$ erg s-1, indicating a decrease of approximately $30\%$ over the four-year period. The bolometric luminosities for the two modes are consistent with each other. However, the Eddington ratios ($\lambda_{\rm Edd}=L_{\rm Bol}/L_{\rm Edd}$) differ significantly due to a factor of $\sim 4$ difference in the median recovered $M_{\rm BH}$ values for the two modes. The corresponding values are listed in Table 2 and illustrated in Fig. 6.
@@ 4.1 The Nature of GSN 069 Emission
One of the key differences between TDEs and standard AGN lies in the radial extent of their respective accretion flows. The characteristic size of the outer disk edge in both systems is set by the origin of the material which is being accreted. In AGN, the accretion flow is sustained by large-scale (circumnuclear) gas transport within the host galaxy (e.g., Storchi-Bergmann & Schnorr-Müller, 2019; Hopkins et al., 2024a, b), and is long-lived, with the accretion flow extending to scales of at least $\mathcal{O}(10^{5}\,R_{g})$.
@@
This large-scale feeding mechanism also gives rise to the additional structures observed in AGN systems (in which the material typically exists in distinct physical states or compositions compared to the optically thick, ionized accretion flow). For example the Broad Line Region (BLR) and the dusty ‘torus’ (e.g., Peterson et al., 2004; Cackett et al., 2021; Hopkins, 2024), typically extend to even larger scales than the disk itself. Our inferred $R_{\rm out}$ confidently rules out the presence of such large-scale accretion flows in GSN 069. Combined with the absence of all other ubiquitous AGN features, as previously discussed, our findings definitively exclude earlier suggestions that GSN 069 could be a ‘true Type 2’ AGN (e.g., Saxton et al., 2011; Miniutti et al., 2013; Shu et al., 2018) [^6] In the AGN unification model, all AGN are intrinsically similar, differing only in luminosity and producing the same broad emission line spectrum. The absence of broad emission lines in Type 2 AGN is explained by the viewing angle relative to the toroidal obscuration surrounding the central engine (Antonucci, 1993). In contrast, ‘true Type 2’ AGN are hypothesized to be long-lived systems that inherently lack a broad line region (BLR). However, the existence of such AGN remains highly debated (Antonucci, 2012), as many initial candidates were later shown either to possess a BLR (e.g., Bianchi et al., 2019) or to be TDEs.. The position of GSN 069’s nuclear optical spectra emission line ratios in the region above the Kewley et al. (2001) line of the BPT diagram (a.k.a the ‘AGN region’) can be explained by the formation of a compact narrow line region produce by the $\gtrsim 10$ year of high ionizing luminosity produced by GSN 069, as shown by Patra et al. (2023), and does not require a long-lived AGN flow.
@@
In a TDE, however, the source of the accreting material is initially deposited very close to the MBH. Specifically, the initial disk size is set (to leading order) by the pericenter radius of the orbit of the star which was disrupted. For a star to be fully disrupted, this is naturally significantly closer to the black hole than the scale of circumnuclear gas feeding in an AGN. This tidal radius can be written as a function of the black hole and disrupted star properties: $R_{T}\approx R_{*}\left(M_{\rm BH}/M_{*}\right)^{1/3}$, with $R_{*}$ and $M_{*}$ being respectively the stellar radius and mass. For typical black hole and stellar parameters, the initial disk size should extend to a few tens to a few hundreds of $R_{g}$. For example, a solar type star disrupted by a $10^{6}M_{\sun}$ MBH would have a tidal radius of $\sim 50R_{g}$. Clearly, TDE disks are expected to be much more compact (initially) than an AGN flow.
@@
This compact disk will then evolve, with material being lost through the inner edge of the flow (i.e., the disk is accreted). The accretion process conserves angular momentum, and as the total mass of the disk drops, some material must therefore be sent to larger radii to conserve the total angular momentum of the flow $(J_{\rm disk}\propto M_{\rm disk}(t)\sqrt{R_{\rm out}(t)}={\rm constant})$. This means that while initially very compact ($R_{\rm out}\sim 10$’s $R_{g}$), as time elapses the disk will expand ($R_{\rm out}\sim 100$-$1000$’s $R_{g}$). Classical disk solutions predict a power-law time-dependence of the disk outer edge ($R_{\rm out}\propto(t/t_{\rm visc})^{3/8}$ for the canonical Cannizzo et al. 1990 model, where $t_{\rm visc}$ is the so-called “viscous” timescale of the flow). As can be seen from this expression, the expansion rate of the disk is strongest at early times $v_{\rm out}\equiv{\rm d}R_{\rm out}/{\rm d}t\sim t^{-5/8}$, meaning the initial expansion is pronounced, followed by a gradual slowing of the expansion rate (physically this is simply because the mass of the disk drops most quickly at early times, so the expansion must be rapid to conserve $J_{\rm disk}$).
@@
An actual inference of the numerical value of $t_{\rm visc}$ for GSN 069’s disk lies beyond the scope of this work, as it requires fully time-dependent modeling of the flow (e.g., Mummery & Balbus, 2020; Mummery et al., 2024a), but will be the subject of future investigations. Nevertheless, since the time of disruption is not well known for GSN 069 our first $R_{\rm out}$ estimates date to at least 4 years after the initial disk formation. As the expansion of the disk is most rapid in these first 4 years, an inferred $R_{\rm out}$ of $\mathcal{O}(10^{3}\,R_{g})$ is certainly a plausible expectation for a TDE disk, at such late times. For example, this is just a factor of ${\rm few}$ larger than the estimated, around similar period ($\sim$ 4 years), for AT2019qiz (Nicholl et al., 2024).
@@ 4.2 Implications for QPE models
For instance, models invoking binary MBH lensing appear to be disfavored due to the observed spectral evolution during eruptions (Miniutti et al., 2023a; Arcodia et al., 2022, 2024; Giustini et al., 2024). Similarly, models that do not involve an accretion disk at all, are also disfavored, as a radiatively efficient disk is required to account for the quiescent (in between the eruptions) emission, now with its optical/UV counterpart also observed for three out of seven QPE sources: GSN 069, eRO-QPE2 (Wevers et al., in prep.), and AT2019qiz (Nicholl et al., 2024).
@@ 4.2.1 Disk Instabilities
One way to stabilize a higher $\lambda_{\rm Edd}$ disk would be to have efficient advective cooling in the innermost regions (the X-ray emitting zones). However, such effects only dominate at $\lambda_{\rm Edd}\gtrsim 0.68$ (Kaur et al., 2023), whereas $\lambda_{\rm Edd}^{2014}<0.68$. Another option would be to ‘fine tune’ $p_{0}^{2014}$ (for some $\alpha>0.7$ value) so that $\lambda_{\rm Edd}^{2014}<\lambda_{\rm min,Edd}^{2014}$ is impose. This, however, would imply that the QPE recurrence time, $T_{\rm QPE}$, is in general a time-dependent quantity, i.e., $T_{\rm QPE}(t)$, and approximately a linear function of $\lambda_{\rm Edd}(t)$, as per Eq. 6. Yet, Miniutti et al. (2023b) analyzed GSN 069’s X-ray light curves across different quiescent $\lambda_{\rm Edd}$ values (e.g., comparing their XMM5 and XMM6 observations), and while they have identified a correlation between $\lambda_{\rm Edd}$ and the intensity of the eruptions, they found no systematic variation in $T_{\rm QPE}$ with changing $\lambda_{\rm Edd}$. This challenges the viability of such fine tuning approach.
@@ 4.2.2 Orbiter/Disk Interactions
Clearly, the fact that the GSN 069 disk was large enough to be intercepted by an orbiting body in 2018 is supportive of the general disk-orbiter collision paradigm. On the contrary, the fact that in 2014 the disk was already large enough to be impacted is difficult to reconcile in this framework. It is crucial to note that a disk more compact than $R_{\rm orb}$ at any of the analyzed epochs is entirely inconsistent with the observed UV spectral shape. Given the measured temperature, such a compact disk would exhibit a UV spectrum at least partially dominated by the Rayleigh-Jeans tail, whereas the observed spectra are dominated by the mid-frequency range shape. Moreover, if the disk were smaller than $R_{\rm orb}$ in 2014 but expanded to exceed $R_{\rm orb}$ by 2018, corresponding to a factor-of-few increase in size, a significant change in the UV spectral shape would be expected (see Figure 1 and 2 in (Guolo & Mummery, 2024)).
One possible interpretation of these results is that disk-orbiter collisions are not the origin of QPEs, and therefore the size of the disk in 2014 and 2018 is irrelevant to the onset (or not) of QPEs. This, trivially, would explain an apparent lack of correlation between disk size and QPE emission. Alternatively, it is reasonable to expect (and this is certainly the case within the framework of e.g., Linial & Metzger, 2023) that more than just the physical extent of the disk determines whether or not QPE emission can be detected. One key part of the model of Linial & Metzger (2023) is that to reach the high measured temperatures of QPE flares ($kT\sim 100$ eV), the gas shock-heated in the collision must be photon-starved, as if it was in thermodynamic equilibrium it would be too cool ($kT\sim 10$ eV) to be detected. Photon starvation describes the state of a gas which is unable to produce enough photons to carry away the luminosity budget of the gas in thermodynamic equilibrium, and is an effect which occurs in supernovae Nakar & Sari (2010), and even black hole accretion flows in extreme regimes (Davis & El-Abd, 2019; Mummery et al., 2024). As photon production by Bremsstrahlung scales with the density of the gas in the disk squared, it is possible that the collisions in 2014 would not have produced photon starved gas, and therefore the flares would be too cool to have been detected. Then, as more material is lost through the inner edge of the disk, the disk density would drop to the point where in 2018 it is possible that the collisions produce gas which is photon starved, and the apparent temperature rises to detectable levels.
@@
In the Lu & Quataert (2023) model the orbiting body transfers mass to the disk through Roche-Lobe overflow, but a collisional model may also add material to the disk through ram pressure induced stripping (Linial & Metzger, 2023; Yao et al., 2024). While our observations do not rule out the collisional model, it does imply that the mass added from collisions must be significantly smaller than the local mass content of the disk at this radius. In particular the decaying temperature from discovery to around 2020 (Miniutti et al., 2023b) means that in this period the total mass of the disk is decreasing and there was no significant mass replenishing, at least until the late-2020 rebrightening.

### v2

@@ 2.2 HST/STIS UV Spectra
Although the signal-to-noise (S/N) ratio is high, no strong broad emission lines typically seen in Seyfert galaxies and quasars (Shull et al., 2012; Stevans et al., 2014) are detected. In this sense, the UV spectrum is mostly ‘featureless’, in comparison with AGN. This absence of lines is unsurprising, as GSN 069 lacks many of the classic AGN features–not only broad optical emission lines (indicating the absence of a mature broad line region, Wevers et al. 2022), but also no hot dust “torus” emission nor hard X-ray corona emission (Miniutti et al., 2019). A comparison to optically selected TDEs may also be warranted: their broad emission lines, usually, used to tell them apart from other optical transients, are associated with their early time ($\leq$ 1 year) ‘UV/optical excess’ emission, whose origin is still debated but is known not to be power by direct disk emission; these lines and this emission component are not present at their later times ($>$ 1 year) when pure disk emission dominates the multi-wavelength TDE emission (Charalampopoulos et al., 2022; van Velzen et al., 2019; Mummery et al., 2024b). If GSN 069 does indeed have a TDE origin, then this early-time UV/optical emission phase was never observed in GSN 069, and all the data presented in this work were taken many years after this initial phase.
@@ 3.1 Fitting Setup and Modeling in the Newtonian Limit
More physically meaningful than $R_{\rm out}/R_{\rm in}$ and $R_{\rm in}^{*}$ would be estimates of $R_{\rm out}$ (in units of the gravitational radius, $R_{\rm g}$) and $R_{\rm in}$. The former would allow for direct comparison with $R_{\rm orb}$, while the latter, when associated with the innermost stable circular orbit (ISCO), is directly correlated with the black hole mass ($M_{\rm BH}$), and would allow for $M_{\rm BH}$ estimates. However, given that in the Newtonian limit of diskSED $i$ and $a$ cannot be marginalized, $R_{\rm out}/R_{\rm g}$ and $M_{\rm BH}$ estimates can only be obtained by assuming ad hoc values (or probability distributions) of $a$ and $i$ (see e.g., Wevers et al., 2025; Guolo & Mummery, 2024). Instead, given the high-quality available data on GSN 069, i.e., two epochs of very high S/N X-ray spectrum and wide UV coverage, estimates on inclination and spin may be obtainable via introduction of relativistic effects into the disk modeling. This is the focus of the next section.
@@ 3.2 Relativistic Disk Modeling
$M_{\rm BH}\,[M_{\odot}]$ $7.5^{+1.0}_{-1.5}\times 10^{6}$ $1.5^{+0.5}_{-0.3}\times 10^{6}$
@@
$L_{\rm Bol}^{\rm 2014}$ [erg s-1] $5.5_{-0.2}^{+0.5}\times 10^{43}$ $3.8_{-0.3}^{+0.3}\times 10^{43}$
$L_{\rm Bol}^{\rm 2018}$ / $L_{\rm Bol}^{\rm 2014}$ $0.69\pm 0.01$ $0.74\pm 0.01$
$\lambda_{\rm Edd}^{\rm 2014}$ $0.05^{+0.01}_{-0.01}$ $0.20^{+0.05}_{-0.05}$
$\lambda_{\rm Edd}^{\rm 2018}/\lambda_{\rm Edd}^{\rm 2014}$ $0.69\pm 0.01$ $0.74\pm 0.01$
@@
Following this approach, we obtain $M_{\rm BH}$ $=5.5^{+1.1}_{-1.5}\times 10^{6}$ $M_{\odot}$ (high spin mode) and $M_{\rm BH}$ $=1.5^{+0.5}_{-0.3}\times 10^{6}$ $M_{\odot}$ (low spin mode). These values are consistent with host-galaxy scaling relations. The nuclear stellar velocity dispersion of GSN 069’s host is $\sigma_{*}=63\pm 4$ km s-1 (Wevers et al., 2022), which translates into $M_{\rm BH}$ estimates of ${\rm few}\times 10^{5}-{\rm few}\times 10^{6}$ $M_{\odot}$ (for the central value only, not adding the intrinsic dispersions) using distinct $M_{\rm BH}$ $-\sigma_{*}$ relations. These relations, however, have systematic $1\sigma$ dispersions of approximately ± ${\pm\,0.5\ \rm dex}$. Thus, even considering the full spin posterior distribution, our $M_{\rm BH}$ estimates are more precise than those derived from host-scaling relations, as our 99% credible interval is (1, 9)×$\times 10^{6}$ $M_{\odot}$, more importantly, our solutions for distinct black hole masses have distinct associated spins and inclination distributions (Table 2).
@@
With the multi-wavelength fitting performed, we can now obtain more precise estimates of the bolometric disk luminosity ($L_{\rm Bol}$) compared to estimates based solely on X-ray data. For the two epochs, we find $L_{\rm Bol}^{2014}=\rm{few}\times 10^{43}$ erg s-1, with slightly deference between the modes, given the anisotropy of the relativistic disk, owing the angle-dependent relativistic effects, a decrease of approximately $30\%$ over the four-year period is observed in both modes. The Eddington ratios ($\lambda_{\rm Edd}=L_{\rm Bol}/L_{\rm Edd}$) however, differ more significantly mainly due to a factor of $\sim 5$ difference in the median recovered $M_{\rm BH}$ values for the two modes. The corresponding values are listed in Table 2 and illustrated in Fig. 6.
@@ 4.1 The Nature of GSN 069 Emission
One of the key differences between TDEs and standard AGN lies in the radial extent of their respective accretion flows. The characteristic size of the outer disk edge in both systems is set by the origin of the material which is being accreted. In AGN, the accretion flow is sustained by large-scale (circumnuclear) gas transport within the host galaxy (e.g., Storchi-Bergmann & Schnorr-Müller, 2019; Angles-Alcazar et al., 2021; Hopkins et al., 2024a, b), and is long-lived, with the accretion flow extending to scales of at least $\mathcal{O}(10^{5}\,R_{g})$.
@@
This large-scale feeding mechanism also gives rise to the additional structures observed in AGN systems (in which the material typically exists in distinct physical states or compositions compared to the optically thick, ionized accretion flow). For example the Broad Line Region (BLR) and the dusty ‘torus’ (e.g., Peterson et al., 2004; Cackett et al., 2021; Hopkins, 2024), typically extend to even larger scales than the disk itself. Our inferred $R_{\rm out}$ confidently rules out the presence of such large-scale accretion flows in GSN 069. Combined with the absence of all other ubiquitous AGN features, as previously discussed, our findings definitively exclude earlier suggestions that GSN 069 could be a ‘true Type 2’ AGN (e.g., Saxton et al., 2011; Miniutti et al., 2013; Shu et al., 2018) [^6] In the AGN unification model, all AGN are intrinsically similar, differing only in luminosity and producing the same broad emission line spectrum. The absence of broad emission lines in Type 2 AGN is explained by the viewing angle relative to the toroidal obscuration surrounding the central engine (Antonucci, 1993). In contrast, ‘true Type 2’ AGN are hypothesized to be long-lived systems that inherently lack a broad line region (BLR). However, the existence of such AGN remains highly debated (Antonucci, 2012), as many initial candidates were later shown either to possess a BLR (e.g., Bianchi et al., 2019) or to be TDEs.. The position of GSN 069’s nuclear optical spectra emission line ratios in the region above the Kewley et al. (2001) line of the BPT diagram (a.k.a the ‘AGN region’) can be explained by the formation of a compact narrow line region produce by the $\gtrsim 10$ year of high ionizing luminosity produced by GSN 069, as shown by e.g., Patra et al. (2023) and Mummery et al. (2025), and does not require a long-lived AGN flow.
@@
In a TDE, however, the source of the accreting material is initially deposited very close to the MBH. Specifically, the initial disk size is set (to leading order) by the pericenter radius of the orbit of the star which was disrupted (Rees, 1988; Cannizzo et al., 1990; Gezari, 2021). For a star to be fully disrupted, this is naturally significantly closer to the black hole than the scale of circumnuclear gas feeding in an AGN. This tidal radius can be written as a function of the black hole and disrupted star properties: $R_{T}\approx R_{*}\left(M_{\rm BH}/M_{*}\right)^{1/3}$, with $R_{*}$ and $M_{*}$ being respectively the stellar radius and mass. For typical black hole and stellar parameters, the initial disk size should extend to a few tens to a few hundreds of $R_{g}$. For example, a solar type star disrupted by a $10^{6}M_{\sun}$ MBH would have a tidal radius of $\sim 50R_{g}$. Clearly, TDE disks are expected to be much more compact (initially) than an AGN flow.
@@
This compact disk will then evolve (e.g., Cannizzo et al., 1990; Mummery & Balbus, 2020), with material being lost through the inner edge of the flow (i.e., the disk is accreted). The accretion process conserves angular momentum, and as the total mass of the disk drops, some material must therefore be sent to larger radii to conserve the total angular momentum of the flow $(J_{\rm disk}\propto M_{\rm disk}(t)\sqrt{R_{\rm out}(t)}={\rm constant})$. This means that while initially very compact ($R_{\rm out}\sim 10$’s $R_{g}$), as time elapses the disk will expand ($R_{\rm out}\sim 100$-$1000$’s $R_{g}$). Classical disk solutions predict a power-law time-dependence of the disk outer edge ($R_{\rm out}\propto(t/t_{\rm visc})^{3/8}$ for the canonical Cannizzo et al. 1990 model, where $t_{\rm visc}$ is the so-called “viscous” timescale of the flow). As can be seen from this expression, the expansion rate of the disk is strongest at early times $v_{\rm out}\equiv{\rm d}R_{\rm out}/{\rm d}t\sim t^{-5/8}$, meaning the initial expansion is pronounced, followed by a gradual slowing of the expansion rate (physically this is simply because the mass of the disk drops most quickly at early times, so the expansion must be rapid to conserve $J_{\rm disk}$).
@@
An actual inference of the numerical value of $t_{\rm visc}$ for GSN 069’s disk lies beyond the scope of this work, as it requires fully time-dependent modeling of the flow (e.g., Mummery & Balbus, 2020; Mummery et al., 2024a), but will be the subject of future investigations. Nevertheless, since the time of disruption is not well known for GSN 069 our first $R_{\rm out}$ estimates date to at least 4 years after the initial disk formation. As the expansion of the disk is most rapid in these first 4 years, an inferred $R_{\rm out}$ of $\mathcal{O}(10^{3}\,R_{g})$ is certainly a plausible expectation for a TDE disk, at such late times. For example, this is just a factor of ${\rm few}$ larger than the estimated for AT2019qiz $\sim$ 4 years after disruption (Nicholl et al., 2024).
@@ 4.2 Implications for QPE models
For instance, models invoking binary MBH lensing appear to be disfavored due to the observed spectral evolution during eruptions (Miniutti et al., 2023a; Arcodia et al., 2022, 2024; Giustini et al., 2024). Similarly, models that do not involve an accretion disk at all, are also disfavored, as a radiatively efficient disk is required to account for the quiescent (in between the eruptions) emission, now with its optical/UV counterpart also observed for three out of seven QPE sources: GSN 069, eRO-QPE2 (Wevers et al., 2025), and AT2019qiz (Nicholl et al., 2024).
@@ 4.2.1 Disk Instabilities
One way to stabilize a higher $\lambda_{\rm Edd}$ disk would be to have efficient advective cooling in the innermost regions (the X-ray emitting zones). However, such effects only dominate at $\lambda_{\rm Edd}\gtrsim 0.68$ (Kaur et al., 2023), whereas $\lambda_{\rm Edd}^{2014}<0.68$. Another option would be to ‘fine tune’ $p_{0}^{2014}$ (for some $\alpha>0.7$ value) so that $\lambda_{\rm Edd}^{2014}<\lambda_{\rm min,Edd}^{2014}$ is imposed. This, however, would imply that the QPE recurrence time, $T_{\rm QPE}$, is in general a time-dependent quantity, i.e., $T_{\rm QPE}(t)$, and approximately a linear function of $\lambda_{\rm Edd}(t)$, as per Eq. 6. Yet, Miniutti et al. (2023b) analyzed GSN 069’s X-ray light curves across different quiescent $\lambda_{\rm Edd}$ values (e.g., comparing their XMM5 and XMM6 observations), and while they have identified a correlation between $\lambda_{\rm Edd}$ and the intensity of the eruptions, they found no systematic variation in $T_{\rm QPE}$ with changing $\lambda_{\rm Edd}$. This challenges the viability of such fine tuning approach.
@@ 4.2.2 Orbiter/Disk Interactions
Clearly, the fact that the GSN 069 disk was large enough to be intercepted by an orbiting body in 2018 is supportive of the general disk-orbiter collision paradigm. On the contrary, the fact that in 2014 the disk was already large enough to be impacted is difficult to reconcile in this framework. It is crucial to note that a disk more compact than $R_{\rm orb}$ at any of the analyzed epochs is entirely inconsistent with the observed UV spectral shape. Given the measured temperature, such a compact disk would exhibit a UV spectrum at least partially dominated by the Rayleigh-Jeans tail, whereas the observed spectra are dominated by the mid-frequency range shape. Moreover, if the disk were smaller than $R_{\rm orb}$ in 2014 but expanded to exceed $R_{\rm orb}$ by 2018, corresponding to a factor-of-few increase in size, a significant change in the UV spectral shape would be expected (see Figure 1 and 2 in Guolo & Mummery, 2024).
One possible interpretation of these results is that disk-orbiter collisions are not the origin of QPEs, and therefore the size of the disk in 2014 and 2018 is irrelevant to the onset (or not) of QPEs. This, trivially, would explain an apparent lack of correlation between disk size and QPE emission. Alternatively, it is reasonable to expect (and this is certainly the case within the framework of e.g., Linial & Metzger, 2023) that more than just the physical extent of the disk determines whether or not QPE emission can be detected. One key part of the model of Linial & Metzger (2023) is that to reach the high measured temperatures of QPE flares ($kT\sim 100$ eV), the gas shock-heated in the collision must be photon-starved, as if it was in thermodynamic equilibrium it would be too cool ($kT\sim 10$ eV) to be detected. Photon starvation describes the state of a gas which is unable to produce enough photons to carry away the luminosity budget of the gas in thermodynamic equilibrium, and is an effect which occurs in supernovae Nakar & Sari (2010), and even black hole accretion flows in extreme regimes (Davis & El-Abd, 2019; Mummery et al., 2024). As photon production by Bremsstrahlung scales with the density of the gas in the disk squared, it is possible that the collisions in 2014 would not have produced photon starved gas, and therefore the flares would be too cool to have been detected. Meaning, that the eruption should till be there, but their lower temperature when in thermodynamic equilibrium ($kT\sim 10$ eV), would make the shocked gas to shine mostly in the unobserved extreme UV wavelengths, and therefore not detectable in the X-rays. Then, as more material is lost through the inner edge of the disk, the disk density would drop to the point where in 2018 it is possible that the collisions produce gas which is photon starved, and the apparent temperature rises to X-ray detectable levels.
@@
In the Lu & Quataert (2023) model the orbiting body transfers mass to the disk through Roche-Lobe overflow, but a collisional model may also add material to the disk through ram pressure induced stripping (Linial & Metzger, 2023; Yao et al., 2024). While our observations do not rule out the collisional model, it does imply that the mass added from collisions must be significantly smaller than the local mass content of the disk at this radius. In particular the decaying temperature from discovery to around 2020 (Miniutti et al., 2023b) means that in this period the inner disk mass is decreasing and there was no significant mass replenishing, at least until the late-2020 rebrightening.